OpenStudy (anonymous):

what is equivalent to log 8 35 rounded to three decimal places

OpenStudy (kinggeorge):

Do you have a calculator that can do logs?

OpenStudy (anonymous):

yes how do i plug that in

OpenStudy (kinggeorge):

Assuming the question is \[\large \log_{8}{35}\]Plug in:\[\Large {\log 35 \over \log 8}\]Where \(\log\) is the log base 10.

OpenStudy (anonymous):

i got 1.70976100.....

OpenStudy (anonymous):

would my answer be 1.710

OpenStudy (anonymous):

??

OpenStudy (kinggeorge):

That is also what I'm getting. Sorry I was gone for a few minutes there.

OpenStudy (anonymous):

its fine! how would i answer thiss? what is equivalent to log 25 125 ?

OpenStudy (kinggeorge):

Here's a nice easy way to see this. Since \[\Large \log_b(a) = x\;\;\;\; \Longrightarrow b^x =a\]So we want some number \(x\) s.t. \[\large 25^x=125\]To solve this, write \(25=5^2\) and \(125=5^3\). Thus, we have \[(5^2)^x =5^{2x}=5^3\]So \(2x=3\). Solving for \(x\) we have \(x=1.5\)

OpenStudy (kinggeorge):

Alternatively, calculate \[\log 125 \over \log25\]

OpenStudy (kinggeorge):

g2g now, good luck.